# Verifying Well-Posedness of Linear PDEs using Convex Optimization

### 01 Apr 2026 v1

@@ III-B The PIE Representation of Linear 1D PDEs
parameterized by $\boldsymbol{A}_{k}\in\mathbb{R}^{n\times n}[s]$ for $k\in\{0,\ldots,d\}$. Note that this parameterization supports a broad class of linear, 1D PDEs, several examples of which are provided in Sec. VI. Given this class of PDEs, we will consider the following definition of well-posedness throughout the paper.
@@
For a given subspace $D\subseteq L_{2}^{n}[a,b]$ and $A:D\to L_{2}^{n}[a,b]$, we say that the PDE defined by $A$ is well-posed if $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$.
Although not immediately apparent from the definition, well-posedness of a given PDE depends not only on the operator $A$, but also on the domain $D$ and the underlying Hilbert space $L_{2}^{n}[a,b]$ on which the system is defined. This distinction is particularly important for PDE models that admit multiple representations in terms of different operators $A$ and domains DD—such as the wave equation in Subsection VI-C—for which well-posedness of one representation does not necessarily imply well-posedness of another. The proposed parameterizations of $D$ (Defn. 2) and $A$ (Eqn. 4) will allow for verification of well-posedness for different such representations. However, well-posedness will only be verified on the Hilbert space $L_{2}^{n}[a,b]$, with extensions to alternative Hilbert spaces left for future study.
@@
$\mathcal{A}:=A\circ\mathcal{T}:=\sum_{k=0}^{d}\textnormal{M}_{\boldsymbol{A}_{k}}\circ\partial_{s}^{k}\circ\mathcal{T},$
for $\textnormal{M}_{\boldsymbol{A}}$ the multiplier operator defined by $\boldsymbol{A}$. It is not difficult to see that for any solution to the PDE, there exists an associated solution to the PIE, and vice versa, establishing equivalence of the two representations [14]. As such, a notion of well-posedness of the PIE can also be defined in terms of the associated PDE as follows.
@@
For $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}\in\mathcal{L}(L_{2}^{n}[a,b])$ bounded linear operators, with $\mathcal{T}$ bijective onto $D$, we say that the PIE defined by $\{\mathcal{T},\mathcal{A}\}$ is well-posed if $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}$ is the infinitesimal generator of a $C_{0}$-semigroup. We refer to the $C_{0}$-semigroup generated by $\{\mathcal{T},\mathcal{A}\}$ as that generated by $A$.
By this definition, well-posedness of a given PIE is equivalent to well-posedness of its associated PDE. The key advantage of the PIE formulation, however, is that both $\mathcal{T}$ and $\mathcal{A}$ are bounded operators on $L_{2}$, and the boundary conditions and regularity constraints are embedded in $\mathcal{T}$. This will allow conditions for well-posedness of the PIE to be formulated in terms of operator inequalities and tested using convex optimization, as we show in the following sections.
@@ IV A Lumer–Phillips Theorem for PIEs
To establish conditions for well-posedness in the PIE representation, in this section, we first propose a PIE-based formulation of the classical Lumer–Phillips generation theorem. This result may be used to verify generation of contraction and quasicontraction semigroups, defined as follows.
@@
For a Hilbert space $H$, a $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $H$ is a quasocontraction semigroup with rate $\omega\in\mathbb{R}$ if $\left\lVert{S(t)}\right\rVert_{H}\leq e^{\omega t}$ for all $t\geq 0$. A contraction semigroup on $H$ is a quasicontraction semigroup with rate $\omega=0$.
Generators of contraction semigroups correspond to PDEs for which the origin is Lyapunov stable, so that the functional $V(\mathbf{u}):=\left\lVert{\mathbf{u}}\right\rVert_{H}^{2}$ is non-increasing along solutions. This requires the generator $A:D\to H$ to be dissipative, i.e., $\left\langle\mathbf{u},A\mathbf{u}\right\rangle\leq 0$ for all $\mathbf{u}\in D$—a condition that is also sufficient for well-posedness of (nonlinear) ODEs. For PDEs, however, an additional condition is required to ensure that $e^{tA}$ is well-defined, as provided by the Lumer–Phillips theorem (Thm. 3.4.2 of [18]).
@@
For a given Hilbert space $H$ and subspace $D\subseteq H$, let $A:D\to H$ be a linear operator. Then $A$ is the infinitesimal generator of a contraction semigroup on $H$ if and only if $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ and the operator $\lambda I-A:D\to H$ is surjective for any (or equivalently all) $\lambda>0$.
Although the Lumer–Phillips theorem only characterizes generators of contraction semigroups, it is well-known that $A$ is the generator of a quasicontraction semigroup with rate $\omega$ if and only if $A-\omega I$ is the generator of a contraction semigroup [17] – making the Lumer–Phillips a powerful tool for verifying generation of quasicontraction semigroups. However, while various techniques have been developed to verify the dissipativity condition of the Lumer–Phillips theorem, testing surjectivity of $\lambda I-A$ for $\lambda>0$ is typically more challenging. In the following subsections, we reformulate both of these conditions in terms of the operators $\{\mathcal{T},\mathcal{A}\}$ defining the associated PIE, using the fact that $\mathcal{T}:L_{2}^{n}\to D$ is bijective to show that both conditions can be more conveniently tested in the PIE representation. In the subsequent subsection, we then use these conditions to established a PIE-based Lumer–Phillips theorem, providing necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup.
Throughout this section, we will consider a general Hilbert space $H$. In the following section, this freedom in the choice of $H$ will be exploited to verify generation of a more general $C_{0}$-semigroup on $L_{2}$, by choosing $H=(L_{2},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ for a coercive and bounded operator $\mathcal{P}$.
@@ IV-A Dissipativity in the PIE Representation
For a Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}:H\to H$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ if and only if $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$.
@@
By Lem. 8, dissipativity of $A=\mathcal{A}\circ\mathcal{T}^{-1}$ can be verified by testing $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$. Here, while AA is defined only on $D$, the operators $\mathcal{T}$ and $\mathcal{A}$ act on all of $H$. Using the inner product $\left\langle\cdot,\cdot\right\rangle_{H}:=\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}}$ for some operator $\mathcal{P}$, this will allow dissipativity to be tested as an operator inequality on $\mathcal{T},\mathcal{A},\mathcal{P}$, as shown in Subsec. V.
@@ IV-B Surjectivity in the PIE Representation
Consider now the second condition of the Lumer–Phillips theorem: surjectivity of $\lambda I-A:D\to H$ for some $\lambda>0$. Although verifying surjectivity of $\lambda I-A$ for general differential operators $A:W_{2}^{d,n}[a,b]\to L_{2}^{n}[a,b]$ is often nontrivial, Cor. 3 shows that for $D\subseteq W_{2}^{d,n}$ parameterized as in Defn. 2, surjectivity of $\partial_{s}^{d}:D\to L_{2}^{n}$ is implied by PIE compatibility of $D$, with an explicit inverse given by $\mathcal{T}$. Using this fact, sujrectivity of $\lambda I-A=(\lambda\mathcal{T}-\mathcal{A})\mathcal{T}^{-1}:D\to L_{2}^{n}$ is then equivalent to surjectivity of $(\lambda\mathcal{T}-\mathcal{A}):L_{2}^{n}\to L_{2}^{n}$, for any $\lambda>0$. Unlike $\lambda I-A$, however, $\lambda\mathcal{T}-\mathcal{A}$ is a bounded operator, allowing surjectivity to be analysed using the open mapping theorem. In particular, the following lemma (a reformulation of Thm. 4.13 in [19]) shows that this property is equivalent to $(\lambda\mathcal{T}-\mathcal{A})^{*}$ being bounded below.
@@
To show that $\textnormal{ran}(\mathcal{B}^{*})$ is closed, let $(\mathcal{B}^{*}\mathbf{x}_{n})$ be a convergent sequence in $\textnormal{ran}(\mathcal{B}^{*})$ with limit $\mathbf{y}$. Then, for all $m,n\in\mathbb{N}$,
@@
Since $(\mathcal{B}\mathbf{x}_{n})$ is Cauchy, also $(\mathbf{x}_{n})$ is Cauchy, and therefore converges to some $\mathbf{x}\in H$. By continuity of $\mathcal{B}^{*}$, this implies $\mathcal{B}^{*}\mathbf{x}=\mathbf{y}$, whence $\textnormal{ran}(\mathcal{B}^{*})$ is closed and $\textnormal{ran}(\mathcal{B})=H$.
@@
$=\sup_{\mathbf{y}\in U}|\left\langle\mathbf{x},\mathcal{B}\mathbf{y}\right\rangle_{H}|\geq\sup_{\mathbf{z}\in\epsilon U}|\left\langle\mathbf{x},\mathbf{z}\right\rangle_{H}|=\epsilon\left\lVert{\mathbf{x}}\right\rVert_{H}.$
By Lemma 9, surjectivity of $\lambda\mathcal{T}-\mathcal{A}:L_{2}^{n}\to L_{2}^{n}$ can be tested by verifying that $\lambda\mathcal{T}^{*}-\mathcal{A}^{*}$ is bounded below. As for the dissipativity constraint, this condition can be posed as on operator inequality and tested using semidefinite programming, as we show in Sec. V.
@@ IV-C A PIE-Based Generation Theorem for Quasicontraction Semigroups
Having shown how both conditions of the Lumer–Phillips theorem can be tested in the PIE representation, we now combine these results to obtain a PIE-based formulation of the Lumer–Phillips theorem. Here, we remark that if $A:D\to H$ generates a $C_{0}$-semigroup on $H$, then for any $\omega\in\mathbb{R}$, $B:=A+\omega I:D\to H$ also generates a $C_{0}$-semigroup, satisfying $e^{tB}=e^{\omega t}e^{tA}$ for all $t\geq 0$ [17]. Using this observation, and applying Lem. 8 and Lem. 9, we obtain the following necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup.
@@
For a given Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}\in\mathcal{L}(H)$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then for any $\omega\in\mathbb{R}$, $A$ is the infinitesimal generator of a quasicontraction semigroup on $H$ with rate $\omega$ if and only if
@@
For sufficiency, suppose both conditions hold. Define $A_{\omega}:=A-\omega I:D\to H$ and $\mathcal{A}_{\omega}:=A_{\omega}\circ\mathcal{T}=\mathcal{A}-\omega\mathcal{T}\in\mathcal{L}(H)$. Then for all $\mathbf{v}\in H$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}_{\omega}\mathbf{v}\right\rangle_{H}=\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}-\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H}\leq 0.$
By Lem. 8, it follows that $\left\langle\mathbf{u},A_{\omega}\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$. In addition, if $\lambda>\omega$ satisfies $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H}$, then for $\lambda_{\omega}:=\lambda-\omega>0$
$\left\lVert{(\lambda_{\omega}\mathcal{T}^{*}-\mathcal{A}_{\omega}^{*})\mathbf{v}}\right\rVert_{H}=\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H}.$
By Lem. 9, it follows that $(\lambda\mathcal{T}-\mathcal{A}_{\omega})\in\mathcal{L}(H)$ is surjective. Since $\mathcal{T}:H\to D$ is bijective, this implies $(\lambda I-A_{\omega}):=(\lambda\mathcal{T}-A_{\omega})\circ\mathcal{T}^{-1}:D\to H$ is surjective. By the Lumer–Phillips theorem (Thm. 7), it follows that $A_{\omega}=\mathcal{A}_{\omega}\circ\mathcal{T}^{-1}$ generates a contraction semigroup on $H$, implying $A=A_{\omega}+\omega I$ generates a quasicontraction semigroup with rate $\omega$ on $H$.
Now, for necessity, suppose $A$ generates a quasicontraction semigroup with rate $\omega$. Then $A_{\omega}:=A-\omega I$ generates a contraction semigroup. Defining $\mathcal{A}_{\omega}:=A_{\omega}\circ\mathcal{T}=\mathcal{A}-\omega\mathcal{T}$, the Lumer-Phillips theorem then implies $\left\langle\mathbf{u},A_{\omega}\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ and $\lambda_{\omega}I-A_{\omega}:D\to H$ is surjective for all $\lambda_{\omega}>0$. By Lem. 8, it follows that $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}_{\omega}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$, whence $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H}$ for all $\mathbf{v}\in H$. In addition, since $\mathcal{T}:D\to H$ is bijective, $\lambda_{\omega}\mathcal{T}-\mathcal{A}_{\omega}:D\to H$ is surjective for all $\lambda_{\omega}>0$. By Lem. 9, it follows that
$\left\lVert{(\lambda_{\omega}\mathcal{T}^{*}-\mathcal{A}_{\omega}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H},$
for all $\lambda_{\omega}>0$. Setting $\lambda=\lambda_{\omega}+\omega>\omega$, the stated conditions hold. ∎
@@
V A PIE-Based Generation Theorem for Arbitrary $C_{0}$-Semigroups

### 02 Apr 2026 v2

@@ III-B The PIE Representation of Linear 1D PDEs
parameterized by $\boldsymbol{A}_{k}\in\mathbb{R}^{n\times n}[s]$ for $k\in\{0,\ldots,d\}$. Note that this parameterization supports a broad class of linear, 1D PDEs, several examples of which are provided in Section VI. Given this class of PDEs, we will consider the following definition of well-posedness.
@@
For a given subspace $D\subseteq W_{2}^{d,n}[a,b]$ and $A:D\to L_{2}^{n}[a,b]$, we say that the PDE defined by $A$ is well-posed if $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$.
Note that well-posedness of a given PDE depends not only on the operator $A$, but also on the domain $D$ and the underlying Hilbert space on which the system is defined. In order to present a unified result, Defn 4 considers only well-posedness on the Hilbert space $L_{2}^{n}$. This is significant for PDE models such as the wave equation in Subsection VI-C – which is not prima facie well-posed on $L_{2}$.
@@
$\mathcal{A}:=A\circ\mathcal{T}:=\sum_{k=0}^{d}\textnormal{M}_{\boldsymbol{A}_{k}}\circ\partial_{s}^{k}\circ\mathcal{T},$ (6)
defining $(\textnormal{M}_{\boldsymbol{A}_{k}}\mathbf{v})(s):=\boldsymbol{A}_{k}(s)\mathbf{v}(s)$. It is not difficult to see that for any solution to the PDE, there exists an associated solution to the PIE, and vice versa, establishing equivalence of the two representations [14]. As such, a notion of well-posedness of the PIE can also be defined in terms of the associated PDE as follows.
@@
For $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}\in\mathcal{L}(L_{2}^{n}[a,b])$ bounded linear operators, with $\mathcal{T}$ bijective, we say that the PIE defined by $\{\mathcal{T},\mathcal{A}\}$ is well-posed if $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}$ is the infinitesimal generator of a $C_{0}$-semigroup. We refer to the $C_{0}$-semigroup generated by $\{\mathcal{T},\mathcal{A}\}$ as that generated by $A$.
By this definition, well-posedness of a given PIE is equivalent to well-posedness of its associated PDE. The key advantage of the PIE formulation, however, is that both $\mathcal{T}$ and $\mathcal{A}$ are bounded operators on $L_{2}^{n}$, and the boundary conditions and regularity constraints are embedded in $\mathcal{T}$. This will allow conditions for well-posedness of the PIE to be formulated in terms of operator inequalities and tested using convex optimization, as we show in the following sections.
@@ IV A Lumer–Phillips Theorem for PIEs
To establish well-posedness conditions in the PIE representation, we first develop a PIE-based formulation of the classical Lumer–Phillips generation theorem. This result enables verification of contraction and quasicontraction semigroups, defined as follows.
@@
For a Hilbert space $H$, a $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $H$ is a quasicontraction semigroup with rate $\omega\in\mathbb{R}$ if $\left\lVert{S(t)}\right\rVert_{H}\leq e^{\omega t}$ for all $t\geq 0$. A contraction semigroup on $H$ is a quasicontraction semigroup with rate $\omega=0$.
Generators of contraction semigroups correspond to PDEs for which the origin is Lyapunov stable, so that the functional $V(\mathbf{u}):=\left\lVert{\mathbf{u}}\right\rVert_{H}^{2}$ is non-increasing along solutions. This requires the generator $A:D\to H$ to be dissipative, i.e., $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$—a condition that is also sufficient for well-posedness of (nonlinear) ODEs. For PDEs, however, an additional condition is required to ensure that $e^{tA}$ is well-defined, as provided by the Lumer–Phillips theorem (Thm. 3.4.2 of [18]).
@@
For a given Hilbert space $H$ and subspace $D\subseteq H$, let $A:D\to H$ be a linear operator. Then $A$ is the infinitesimal generator of a contraction semigroup on $H$ if and only if $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ and the operator $\lambda I-A:D\to H$ is surjective for some (equivalently, all) $\lambda>0$.
Although the Lumer–Phillips theorem characterizes generators of contraction semigroups, it is well known that $A$ generates a contraction semigroup if and only if $A+\omega I$ generates a quasicontraction semigroup with rate $\omega$ [19], allowing direct application of the theorem to this broader class of generators. However, testing the conditions of the Lumer–Phillips – in particular the surjectivity of $\lambda I-A$ for $\lambda>0$ – still requires substantial analysis.
In the following subsections, we reformulate the conditions of the Lumer–Phillips theorem in terms of the associated PIE, defined by $\{\mathcal{T},\mathcal{A}\}$, using the bijectivity of $\mathcal{T}:L_{2}^{n}\to D$ to show that these conditions can be tested more conveniently in the PIE representation. In the subsequent subsection, we then establish a PIE-based Lumer–Phillips theorem, providing necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup. These results are stated for a general Hilbert space $H$, and will be applied to $H=L_{2,\mathcal{P}}^{n}$ in the next section, to verify generation of more general, non-quasicontraction semigroups on $L_{2}^{n}$.
@@ IV-A Dissipativity in the PIE Representation
For a Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}:H\to H$ be bounded linear operators, with $\mathcal{T}$ bijective. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ if and only if $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$.
@@
By Lem. 8, dissipativity of $A=\mathcal{A}\circ\mathcal{T}^{-1}$ can be verified by testing $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$. Here, while AA is defined only on $D$, the operators $\mathcal{T}$ and $\mathcal{A}$ act on all of $H$. Considering $H=L_{2,\mathcal{P}}^{n}$, this will allow dissipativity to be tested as an operator inequality on $\mathcal{T}$, $\mathcal{A}$, and $\mathcal{P}$, as shown in Section V.
@@ IV-B Surjectivity in the PIE Representation
Consider now the second condition of the Lumer–Phillips theorem: surjectivity of $\lambda I-A:D\to H$ for some $\lambda>0$. Although verifying surjectivity of $\lambda I-A$ for general differential operators $A:D\to H$ is nontrivial, Cor. 3 shows that for $H=L_{2}^{n}$ and $D\subseteq W_{2}^{d,n}$ parameterized as in Defn. 2, surjectivity of $\partial_{s}^{d}:D\to H$ is implied by PIE compatibility of $D$, with an explicit inverse given by $\mathcal{T}$. Using this fact, surjectivity of $\lambda I-A=(\lambda\mathcal{T}-\mathcal{A})\mathcal{T}^{-1}:D\to H$ is then equivalent to surjectivity of $(\lambda\mathcal{T}-\mathcal{A}):H\to H$, for any $\lambda>0$. Unlike $\lambda I-A$, however, $\lambda\mathcal{T}-\mathcal{A}$ is a bounded operator, allowing surjectivity to be analysed using the open mapping theorem. In particular, the following lemma (a reformulation of Thm. 4.13 in [20]) shows that this property is equivalent to $(\lambda\mathcal{T}-\mathcal{A})^{*}$ being bounded below.
@@
To show that $\textnormal{ran}(\mathcal{B}^{*})$ is closed, let $(\mathbf{y}_{n})_{n\in\mathbb{N}}$ be a convergent sequence in $\textnormal{ran}(\mathcal{B}^{*})$ with limit $\mathbf{y}$. Then there exists $(\mathbf{x}_{n})_{n\in\mathbb{N}}$ such that $\mathbf{y}_{n}=\mathcal{B}^{*}\mathbf{x}_{n}$. Now $(\mathbf{x}_{n})_{n\in\mathbb{N}}$ is Cauchy, since for all $m,n\in\mathbb{N}$,
@@
and since $(\mathcal{B}^{*}\mathbf{x}_{n})_{n\in\mathbb{N}}$ is convergent and therefore Cauchy. Since $H$ is a Hilbert space, it follows that $(\mathbf{x}_{n})_{n\in\mathbb{N}}$ converges to some $\mathbf{x}\in H$. By continuity of $\mathcal{B}^{*}$ this implies $\mathcal{B}^{*}\mathbf{x}=\mathbf{y}$, whence $\textnormal{ran}(\mathcal{B}^{*})$ is closed and thus $\textnormal{ran}(\mathcal{B})=H$.
@@
$\qquad=\sup_{\mathbf{y}\in U}|\left\langle\mathbf{x},\mathcal{B}\mathbf{y}\right\rangle_{H}|\geq\sup_{\mathbf{z}\in\epsilon U}|\left\langle\mathbf{x},\mathbf{z}\right\rangle_{H}|=\epsilon\left\lVert{\mathbf{x}}\right\rVert_{H}.$
By Lemma 9, surjectivity of $\lambda\mathcal{T}-\mathcal{A}\in\mathcal{L}(H)$ can be tested by verifying that $\lambda\mathcal{T}^{*}-\mathcal{A}^{*}$ is bounded below. As for the dissipativity constraint, this condition can be posed as an operator inequality and tested using semidefinite programming, as we show in Section V.
@@ IV-C A PIE-Based Generation Theorem for Quasicontraction Semigroups
Having shown how both conditions of the Lumer–Phillips theorem can be expressed using the PIE representation, $\{\mathcal{T},\mathcal{A}\}$, we now combine these results to obtain a PIE-based formulation of the Lumer–Phillips theorem. Here, we remark that if $A:D\to H$ generates a $C_{0}$-semigroup on $H$, then for any $\omega\in\mathbb{R}$, $B:=A+\omega I:D\to H$ also generates a $C_{0}$-semigroup, satisfying $e^{tB}=e^{\omega t}e^{tA}$ for all $t\geq 0$ [19]. Using this observation, and applying Lem. 8 and Lem. 9, we obtain the following necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup.
@@
For a given Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}\in\mathcal{L}(H)$ be bounded linear operators, with $\mathcal{T}$ bijective. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then for any $\omega\in\mathbb{R}$, $A$ is the infinitesimal generator of a quasicontraction semigroup with rate $\omega$ on $H$ if and only if
@@
Define $A_{\omega}:=A-\omega I:D\to H$ and $\mathcal{A}_{\omega}:=A_{\omega}\circ\mathcal{T}=\mathcal{A}-\omega\mathcal{T}\in\mathcal{L}(H)$. Then for all $\mathbf{v}\in H$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}_{\omega}\mathbf{v}\right\rangle_{H}=\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}-\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H},$ (*)
and, for any $\lambda>\omega$ and $\lambda_{\omega}:=\lambda-\omega>0$,
$\left\lVert{(\lambda_{\omega}\mathcal{T}^{*}-\mathcal{A}_{\omega}^{*})\mathbf{v}}\right\rVert_{H}=\left\lVert{(\lambda\mathcal{T}-\mathcal{A})^{*}\mathbf{v}}\right\rVert_{H}.$ ($\star$)
[…]
@@
V A PIE-Based Well-Posedness Test for PDEs